### v1

# A conjecture of Radu and Sellers on congruences modulo powers of 2 for broken 3-diamond partitions

Abstract.
In 2007, Andrews and Paule introduced the family of functions $\Delta_{k}(n)$, which enumerate the number of broken $k$-diamond partitions for a fixed positive integer $k$. In 2013, Radu and Sellers completely characterized the parity of $\Delta_{3}(8n+r)$ for some $r$ and gave a conjecture on congruences modulo powers of 2 for broken 3-diamond partitions. We use an unconventional $U$-sequence to solve the revised conjecture proposed by Radu and Sellers.
Key words and phrases:
Broken $k$-diamonds; Congruences; Modular forms; Partitions
2010 Mathematics Subject Classification:
11P83, 05A17
1. Introduction
A partition of a positive integer $n$ is defined as a sequence of positive integers in non-increasing order that sum to $n$. Denoted by $p(n)$, it represents the count of such partitions for $n$. The following distinguished congruences were discovered by Ramanujan in [13]:
$p(5n+4)\equiv 0$ $\pmod{5},$
$p(7n+5)\equiv 0$ $\pmod{7},$
$p(11n+6)\equiv 0$ $\pmod{11},$
for every non-negative integer $n$. In fact, for $\alpha\geq 1$
(1.1) $p(5^{\alpha}n+\delta_{5,\alpha})$ $\equiv 0\pmod{5^{\alpha}},$
(1.2) $p(7^{\alpha}n+\delta_{7,\alpha})$ $\equiv 0\pmod{7^{[\frac{\alpha+2}{2}]}},$
(1.3) $p(11^{\alpha}n+\delta_{11,\alpha})$ $\equiv 0\pmod{11^{\alpha}},$
where $\delta_{t,\alpha}$ is the reciprocal modulo $t^{\alpha}$ of 24. (1.1) and (1.2) were first proved by Watson [16] in 1938. For an elementary of (1.1) see Hirschhorn and Hunt [8] and for an elementary proof of (1.2) see Garvan [6]. (1.3) was proved by Atkin [3] in 1967.
In 2007, Andrews and Paule [2] introduced the family of functions $\Delta_{k}(n)$ which enumerate the number of broken $k$-diamond partitions for a fixed positive integer $k$. These are constructed in such a way that the generating functions of their counting sequences $\{\Delta_{k}(n)\}_{n\geq 0}$ are closely related to modular forms. Namely,
$\sum_{n=0}^{\infty}\Delta_{k}(n)q^{n}$ $=\prod_{n=1}^{\infty}\frac{(1-q^{2n})(1-q^{(2k+1)n})}{(1-q^{n})^{3}(1-q^{(4k+2)n})}$
$=q^{(k+1)/12}\frac{\eta(2\tau)\eta((2k+1)\tau)}{\eta(\tau)^{3}\eta((4k+2)\tau)},~{}k\geq 1,$
where we recall the Dedekind eta function
$\eta(\tau)=q^{\frac{1}{24}}\prod_{n=1}^{\infty}(1-q^{n})~{}~{}(q:=e^{2\pi i\tau}).$
For $k=1$, Andrews and Paule proved that for all $n\geq 0$, $\Delta_{1}(2n+1)\equiv 0\pmod{3}$. Hirschhorn and Sellers [9] considered cases of $k=1$ and $k=2$ and gave the parity results below: for all $n\geq 0$,
$\Delta_{1}(4n+2)\equiv 0\pmod{2},$
$\Delta_{1}(4n+3)\equiv 0\pmod{2},$
$\Delta_{2}(10n+2)\equiv 0\pmod{2},$
$\Delta_{2}(10n+6)\equiv 0\pmod{2}.$
Besides, Chan [5] provided a different proof of the parity results for $\Delta_{2}$ mentioned above as well as a number of congruences modulo powers of 5. Paule and Radu [10] also proved a number of congruences modulo 5 for broken 2-diamond partitions, and they also shared conjectures related to broken 3-diamond partitions modulo 7 and broken 5-diamond partitions modulo 11, two of them proved by Xiong [15].
In 2013, Radu and Sellers [12] completely characterized the parity of $\Delta_{3}(8n+r)$ for some $r$ and gave the following conjecture modulo powers of $2$.
Conjecture 1.1.
Let
$\lambda_{\alpha}=\left\{\begin{array}[]{ll}\frac{2^{\alpha+1}+1}{3},{}&\text{if}~{}\alpha~{}\text{is even,}\\ \frac{2^{\alpha}+1}{3},{}&\text{if}~{}\alpha~{}\text{is odd.}\end{array}\right.$
Then for all $\alpha\geq 1$ and $n\geq 0$,
(1.4) $\Delta_{3}(\lambda_{\alpha})\Delta_{3}(2^{\alpha+2}n+\lambda_{\alpha+2})\equiv\Delta_{3}(\lambda_{\alpha+2})\Delta_{3}(2^{\alpha}n+\lambda_{\alpha})\pmod{2^{\alpha}}$
and
(1.5) $\Delta_{3}(\lambda_{\alpha})\equiv 1\pmod{2}.$
They also proved the case for $\alpha=1$. Later, Xia [14] demonstrated that (1.4) holds for $\alpha=2$ and (1.5) holds for all $\alpha\geq 1$. But (1.4) is incorrect in general. For example let $\alpha=3$ and $n=1$, we have
$\Delta_{3}(\lambda_{3})=\Delta_{3}(3)=19,$
$\Delta_{3}(2^{5}+\lambda_{5})=\Delta_{3}(43)=236491535,$
$\Delta_{3}(\lambda_{5})=\Delta_{3}(2^{3}+\lambda_{3})=\Delta_{3}(11)=2653,$
$\Delta_{3}(\lambda_{3})\Delta_{3}(2^{5}+\lambda_{5})-\Delta_{3}(\lambda_{5})\Delta_{3}(2^{3}+\lambda_{3})=2^{2}\cdot 1121575189\not\equiv 0\pmod{2^{3}},$
which indicates that (1.4) does not hold for $\alpha=3$. The correct version of this conjecture, which we prove, is given below in Theorem 4.9.
In this paper, we prove this conjecture by following the method in [11]. We associate the family of congruences with a sequence of functions defined on the classical modular curve $X_{0}(28)$ (a curve of genus 2 and cusp count 6). Let
$F(q):=q^{\frac{1}{3}}\frac{\eta(2\tau)\eta(7\tau)}{\eta(\tau)^{3}\eta(14\tau)}=\sum_{n=0}^{\infty}\Delta_{3}(n)q^{n}.$
We mention that we use an unconventional $U$-sequence because $F(q)/F(q^{4})$ is not a modular function, and there does not exist any $U$-sequence of functions defined on the classical modular curve $X_{0}(14)$ (a curve of genus 1 and cusp count 4) associated with the family of congruences.
The paper is organized as follows. In Section 2 we introduce the necessary notations and definitions. In Section 3 we derive modular equations on $\Gamma_{0}(28)$. Section 4 is dedicated to the proof of revised Conjecture 1.1.
2. Preliminaries
In this paper, we use the following conventions: =$\mathbb{N}^{\ast}=\{1,2,\cdots\}$ denotes the set of positive integers. The complex upper half plane is denoted by :=$\mathbb{H}:=\{\tau\in\mathbb{C}:\Im(\tau)\textgreater 0\}$. We will also use the short hand notation for the Dedekind eta function:
$\eta_{n}(\tau):=\eta(n\tau),~{}~{}n\in\mathbb{Z},~{}\tau\in\mathbb{H}.$
Throughout this paper, for $x\in\mathbb{R}$ the symbol $\lfloor x\rfloor$ denotes the largest integer less than or equal to $x$, and $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$.
Let $f=\sum_{n\in\mathbb{Z}}a_{n}q^{n}$, $f\neq 0$, be such that $a_{n}=0$ for almost all $n\textless 0$. We define the order of $f$ (with respect to $q$) as the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{q}(f)$. For instance, let $F=f\circ t=\sum_{n\in\mathbb{Z}}a_{n}t^{n}$ where $t=\sum_{n\geq 1}b_{n}q^{n}$, then the $t$-order of $F$ is defined to be the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{t}(F)$.
Definition 2.1.
For $f:\mathbb{H}\rightarrow\mathbb{C}$ and $m\in\mathbb{N}^{\ast}$ we define $U_{m}(f):\mathbb{H}\rightarrow\mathbb{C}$ by
$U_{m}(f)(\tau):=\frac{1}{m}\sum_{\lambda=0}^{m-1}f(\frac{\tau+\lambda}{m}),~{}\tau\in\mathbb{H}.$
$U_{m}$ is linear (over $\mathbb{C}$); in addition, it is easy to verify that
$U_{mn}=U_{m}\circ U_{n}=U_{n}\circ U_{m},~{}m,n\in\mathbb{N}^{\ast}.$
The operators $U_{m}$, introduced by Atkin and Lehner [4], are closely related to Hecke operators. They typically arise in the context of partition congruences [1] mostly because of the property: if
$f(\tau)=\sum_{n=-\infty}^{\infty}f_{n}q^{n}~{}~{}(q=e^{2\pi i\tau}),$
then
$U_{m}(f)(\tau)=\sum_{n=-\infty}^{\infty}f_{mn}q^{n}.$
Definition 2.2.
For $f:\mathbb{H}\rightarrow\mathbb{C}$ we define $U^{(0)}(f),U^{(1)}(f):\mathbb{H}\rightarrow\mathbb{C}$ by $U^{(0)}(f):=U_{2}(Af)$ and $U^{(1)}(f):=U_{2}(Bf)$, where $A:=\frac{\eta_{8}^{4}\eta_{2}^{2}}{\eta_{4}^{2}\eta_{1}^{4}}$ and $B:=\frac{\eta_{1}^{4}\eta_{2}^{2}\eta_{8}^{4}}{\eta_{4}^{10}}$.
Definition 2.3.
We define the $U$-sequence $(L_{\alpha})_{\alpha\geq 0}$ by $L_{0}:=\frac{\eta_{7}\eta_{4}^{4}\eta_{1}}{\eta_{14}\eta_{2}^{5}}$ and for $\alpha\geq 1$:
$L_{2\alpha-1}:=U^{(0)}(L_{2\alpha-2})~{}~{}~{}~{}~{}\text{and}~{}~{}~{}L_{2\alpha}:=U^{(1)}(L_{2\alpha-1}).$
The proof of the following lemma is completely analogous to [3, p. 23] and we omit it.
Lemma 2.4.
For $\alpha\in\mathbb{N}^{\ast}$, we have
$L_{2\alpha-1}=q\prod_{n=1}^{\infty}\frac{(1-q^{4n})^{4}(1-q^{2n})^{2}}{(1-q^{n})^{4}}\sum_{n=0}^{\infty}\Delta_{3}(2^{2\alpha-1}n+\lambda_{2\alpha-1})q^{n},$
$L_{2\alpha}=q\prod_{n=1}^{\infty}\frac{(1-q^{n})^{4}(1-q^{4n})^{4}}{(1-q^{2n})^{6}}\sum_{n=0}^{\infty}\Delta_{3}(2^{2\alpha}n+\lambda_{2\alpha})q^{n}.$
Definition 2.5.
Let $t,p,p_{0},y_{0},p_{1}$ be functions defined on $\mathbb{H}$ as follows:
$t:=\frac{\eta_{28}\eta_{4}}{\eta_{7}\eta_{1}},~{}~{}~{}p:=\frac{1}{2}(1+\frac{7\eta_{14}^{7}\eta_{4}^{5}\eta_{1}^{6}}{\eta_{28}^{3}\eta_{7}^{2}\eta_{2}^{13}}),$
$p_{0}:=\frac{\eta_{7}\eta_{4}^{7}}{\eta_{28}\eta_{1}^{7}},~{}~{}~{}y_{0}:=\frac{7\eta_{14}^{5}\eta_{4}^{5}\eta_{2}}{\eta_{28}^{3}\eta_{1}^{8}},~{}~{}~{}p_{1}:=\frac{1}{4}(p+2pt+t)+t^{2}.$
3. Modular equation
Our proof of Radu and Sellers’ conjecture relies on the identities in the Appendix. All these identities can be proved using modular function theory. Regarding to the valence formula in the modular forms, Garvan has written a MAPLE package called ETA which can prove Group I, Group III, and Group IV in the Appendix automatically. See
(3.1) $https://qseries.org/fgarvan/qmaple/ETA/$
The tutorial for this see [7]. The identities in Group II of the Appendix cannot be proved directly since $p_{1}$ is not an eta-quotient. However, we can use the ETA MAPLE package to prove that
(3.2) $U^{(0)}(t^{3})=$ $p_{0}(256t^{7}+640t^{6}+864t^{5}+656t^{4}+276t^{3}+64t^{2}+7t)$
$-y_{0}(32t^{5}+48t^{4}+28t^{3}+8t^{2}+t);$
(3.3) $U^{(0)}(t^{4})=$ $p_{0}(1024t^{9}+3072t^{8}+4864t^{7}+4672t^{6}+2832t^{5}+1104t^{4}+260t^{3}+28t^{2})$
$-y_{0}(128t^{7}+256t^{6}+224t^{5}+112t^{4}+32t^{3}+4t^{2});$
(3.4) $U^{(0)}(yt)=$ $-p_{0}(14t+7)+y_{0}(2t+1);$
(3.5) $U^{(0)}(yt^{2})=$ $-p_{0}(56t^{3}+84t^{2}+28t)+y_{0}(8t^{2}+4t);$
(3.6) $U^{(0)}(yt^{3})=$ $-p_{0}(224t^{5}+448t^{4}+308t^{3}+84t^{2}+7t)+y_{0}(32t^{4}+36t^{3}+12t^{2}+t);$
where $y:=2p-1$ be an eta-quotient. Then we deduce the identities in Group II using (3.2)–(3.6) along with the identities in Group I since $p_{1}=\frac{1}{8}+\frac{1}{2}t+t^{2}+\frac{1}{8}y+\frac{1}{4}yt$.
Definition 3.1.
With $t=t(\tau)$ as in Definition 2.5 we define:
$a_{0}(t)=2t^{2}+t,~{}~{}~{}a_{1}(t)=-4t^{2}-2t.$
We define $s:\{0,1\}\times\{1,2\}\rightarrow\mathbb{Z}$ to be the unique function satisfying
(3.7) $a_{j}(t)=\sum_{l=1}^{2}s(j,l)2^{\lfloor\frac{2l+j-1}{2}\rfloor}t^{l}.$
Lemma 3.2.
For $\lambda=0$ or $1$, let
$t_{\lambda}(\tau):=t(\frac{\tau+\lambda}{2}),~{}~{}\tau\in\mathbb{H}.$
Then in the polynomial ring $\mathbb{C}(t)[X]$:
$X^{2}+a_{1}(t)X+a_{0}(t)=(X-t_{0})(X-t_{1}).$
Proof.
First we prove
$a_{0}(t)=\frac{\eta_{2}^{3}\eta_{14}^{3}}{\eta_{1}^{3}\eta_{7}^{3}}=2t^{2}+t.$
We have
$a_{0}(t)=$ $t_{0}t_{1}=\prod_{\lambda=0}^{1}q^{1/2}\prod_{n=1}^{\infty}\frac{(1-q^{14n})(1-q^{2n})}{(1-(-1)^{\lambda n}q^{7n/2})(1-(-1)^{\lambda n}q^{n/2})}$
$=$ $q\prod_{n=1}^{\infty}(1-q^{14n})^{2}(1-q^{2n})^{2}\prod_{n=1}^{\infty}\frac{1}{(1-q^{2n+1})(1-q^{14n+7})}\prod_{n=1}^{\infty}\frac{1}{(1-q^{n})^{2}(1-q^{7n})^{2}}$
$=$ $q\prod_{n=1}^{\infty}\frac{(1-q^{14n})^{3}(1-q^{2n})^{3}}{(1-q^{7n})^{3}(1-q^{n})^{3}}=2t^{2}+t,$
where the last equation was verified by Garvan’s MAPLE package ETA (3.1).
For $a_{1}(t)$,
$a_{1}(t)=$ $-t_{0}-t_{1}=-2U_{2}(t)=-4t^{2}-2t,$
where the last equation was again verified by the ETA MAPLE package. ∎
Lemma 3.3.
For $u:\mathbb{H}\rightarrow\mathbb{C}$ and $j\in\mathbb{Z}$:
$U_{2}(ut^{j})=-a_{0}(t)U_{2}(ut^{j-2})-a_{1}(t)U_{2}(ut^{j-1}).$
Proof.
For $\lambda\in\{0,1\}$, Lemma 3.2 implies
$t_{\lambda}^{2}+a_{1}(t)t_{\lambda}+a_{0}(t)=0.$
Multiplying both sides with $u_{\lambda}t_{\lambda}^{j-2}$ where $u_{\lambda}(\tau):=u((\tau+\lambda)/2)$ gives
$u_{\lambda}t_{\lambda}^{j}+a_{1}(t)u_{\lambda}t_{\lambda}^{j-1}+a_{0}(t)u_{\lambda}t_{\lambda}^{j-2}=0.$
Summing both sides over all $\lambda$ from $\{0,1\}$ completes the proof of the lemma. ∎
According to Lemma 3.3, we can calculate $U^{(0)}(t^{k})$, $U^{(0)}(p_{1}t^{k})$, $U^{(1)}(y_{0}t^{k})$, and $U^{(1)}(p_{0}t^{k})$ using the identities in Appendix for all $k\geq 1$ recursively. For example
(3.8) $U^{(1)}(y_{0}t^{3})=$ $(4t^{2}+2t)U^{(1)}(y_{0}t^{2})-(2t^{2}+t)U^{(1)}(y_{0}t)$
$=$ $p_{1}(2^{17}t^{13}+2^{19}t^{12}+125\cdot 2^{13}t^{11}+315\cdot 2^{12}t^{10}+2261\cdot 2^{9}t^{9}+1505\cdot 2^{9}t^{8}$
$+6031\cdot 2^{6}t^{7}+1131\cdot 2^{7}t^{6}+4953\cdot 2^{3}t^{5}+1875\cdot 2^{2}t^{4}+441\cdot 2t^{3}+49t^{2})$
$-2^{17}t^{15}-9\cdot 2^{16}t^{14}-145\cdot 2^{13}t^{13}-93\cdot 2^{14}t^{12}-2693\cdot 2^{9}t^{11}-3591\cdot 2^{8}t^{10}$
$-7165\cdot 2^{6}t^{9}-5305\cdot 2^{5}t^{8}-5675\cdot 2^{3}t^{7}-1035\cdot 2^{3}t^{6}-463\cdot 2t^{5}-13\cdot 2^{2}t^{4}.$
4. Radu and Sellers’ conjecture
Lemma 4.1.
Let $v_{1},v_{2},u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $k=l~{}~{}\text{or}~{}~{}l+1$ there exist polynomials $p_{k}^{(1)}(t),p_{k}^{(2)}(t)\in\mathbb{Z}[t]$ such that
(4.1) $U_{2}(ut^{k})=v_{1}p_{k}^{(1)}(t)+v_{2}p_{k}^{(2)}(t)$
and
(4.2) $ord_{t}(p_{k}^{(i)}(t))\geq\lceil\frac{k+s_{i}}{2}\rceil,i\in\{1,2\},$
for some fixed integers $s_{1}$ and $s_{2}$. Then there exist families of polynomials $p_{k}^{(1)}(t),p_{k}^{(2)}(t)\in\mathbb{Z}[t]$, integer $k\geq l$, such that (4.1) and (4.2) hold for integer $k\geq l$.
Proof.
Let $N\textgreater l+1$ be an integer and assume by induction that there are families of polynomials $p_{k}^{(i)}(t),i\in\{1,2\}$, such that (4.1) and (4.2) hold for $l\leq k\leq N-1$. Suppose
$p_{k}^{(i)}(t)=\sum_{n\geq\lceil\frac{k+s_{i}}{2}\rceil}c_{i}(k,n)t^{n},l\leq k\leq N-1,$
with integers $c_{i}(k,n)$. Apply Lemma 3.3 we get:
$U_{2}(ut^{N})=$ $-\sum_{j=0}^{1}a_{j}(t)U_{2}(ut^{N+j-2})$
$=$ $-\sum_{j=0}^{1}a_{j}(t)\sum_{i=1}^{2}v_{i}\sum_{n\geq\lceil\frac{N+j-2+s_{i}}{2}\rceil}c_{i}(N+j-2,n)t^{n}$
$=$ $-\sum_{i=1}^{2}v_{i}\sum_{j=0}^{1}a_{j}(t)t^{-1}\sum_{n\geq\lceil\frac{N+j+s_{i}}{2}\rceil}c_{i}(N+j-2,n-1)t^{n}.$
Recalling the fact that $a_{j}(t)t^{-1}$ for $0\leq j\leq 1$ is a polynomial in $t$, this determines polynomials $p_{N}^{(i)}(t)$ with the desired properties. ∎
Lemma 4.2.
Let $v_{1},v_{2},u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $k=l~{}~{}\text{or}~{}~{}l+1$ there exist polynomials $p_{k}^{(i)}(t)\in\mathbb{Z}[t]$, $i\in\{1,2\}$, such that
(4.3) $U_{2}(ut^{k})=v_{1}p_{k}^{(1)}(t)+v_{2}p_{k}^{(2)}(t)$
where
(4.4) $p_{k}^{(i)}(t)=\sum_{n}c_{i}(k,n)2^{\lfloor\frac{2n-k+\gamma_{i}}{2}\rfloor}t^{n},$
with integers $\gamma_{i}$ and $c_{i}(k,n)$. Then there exist families of polynomials $p_{k}^{(i)}(t)\in\mathbb{Z}[t]$, integer $k\geq l$, of the form (4.4) for which property (4.3) holds for integer $k\geq l$.
Proof.
Suppose for an integer $N\textgreater l+1$ there are families of polynomials $p_{k}^{(i)}(t)$, $i\in\{1,2\}$, of the form (4.4) satisfying property (4.3) for $l\leq k\leq N-1$. We proceed by mathematical induction on $N$. Applying Lemma 3.3 and using the induction base (4.3) and (4.4) we attain:
$U_{2}(ut^{N})=-\sum_{j=0}^{1}a_{j}(t)\sum_{i=1}^{2}v_{i}\sum_{n}c_{i}(N+j-2,n)2^{\lfloor{\frac{2n-(N+j-2)+\gamma_{i}}{2}}\rfloor}t^{n}.$
Utilizing (3.7) this rewrites into:
$U_{2}(ut^{N})$ $=-\sum_{j=0}^{1}\sum_{l=1}^{2}s(j,l)2^{\lfloor{\frac{2l+j-1}{2}}\rfloor}t^{l}\sum_{i=1}^{2}v_{i}\sum_{n}c_{i}(N+j-2,n)2^{\lfloor{\frac{2n-(N+j-2)+\gamma_{i}}{2}}\rfloor}t^{n}$
$=-\sum_{i=1}^{2}v_{i}\sum_{j=0}^{1}\sum_{l=1}^{2}\sum_{n}s(j,l)c_{i}(N+j-2,n-l)2^{\lfloor{\frac{2(n-l)-(N+j-2)+\gamma_{i}}{2}}\rfloor+\lfloor{\frac{2l+j-1}{2}}\rfloor}t^{n}.$
The induction step is completed by simplifying the exponent of $2$ as follows:
$\lfloor{\frac{2(n-l)-(N+j-2)+\gamma_{i}}{2}}\rfloor+\lfloor{\frac{2l+j-1}{2}}\rfloor$
$\geq\lfloor{\frac{2(n-l)-(N+j-2)+\gamma_{i}+2l+j-2}{2}}\rfloor$
$=\lfloor{\frac{2n-N+\gamma_{i}}{2}}\rfloor.$
Thus for $N\geq l$, $U_{2}(ut^{N})$ has the desired expression. ∎
Definition 4.3.
A map $a:\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete function if it has finite support. A map $a:\mathbb{Z}\times\mathbb{Z}\rightarrow\mathbb{Z}$ is called discrete array if for each $i\in\mathbb{Z}$ the map $a(i,-):\mathbb{Z}\rightarrow\mathbb{Z},j\mapsto a(i,j)$, has finite support.
Lemma 4.4.
Given $U^{(0)}(f)$, $U^{(1)}(f)$ as in Definition 2.2, $p_{0},p_{1}$, and $y_{0}$ as in Definition 2.5. Then there exist discrete arrays $a_{i}$, $b_{i}$, $c_{i}$, and $d_{i}$, $i\in\{0,1\}$, such that the following relations hold for all $k\in\mathbb{N^{\ast}}$:
(4.5) $U^{(0)}(t^{k})=p_{0}\sum_{n\geq\lceil\frac{k-1}{2}\rceil}a_{0}(k,n)2^{\lfloor\frac{2n-k+2}{2}\rfloor}t^{n}+y_{0}\sum_{n\geq\lceil\frac{k-1}{2}\rceil}a_{1}(k,n)2^{\lfloor\frac{2n-k+2}{2}\rfloor}t^{n},$
(4.6) $U^{(0)}(p_{1}t^{k})=p_{0}\sum_{n\geq\lceil\frac{k+1}{2}\rceil}b_{0}(k,n)2^{\lfloor\frac{2n-k+1}{2}\rfloor}t^{n}+y_{0}\sum_{n\geq\lceil\frac{k+1}{2}\rceil}b_{1}(k,n)2^{\lfloor\frac{2n-k-1}{2}\rfloor}t^{n},$
(4.7) $U^{(1)}(p_{0}t^{k})=p_{1}\sum_{n\geq\lceil\frac{k}{2}\rceil}c_{0}(k,n)2^{\lfloor\frac{2n-k+1}{2}\rfloor}t^{n}+\sum_{n\geq\lceil\frac{k+4}{2}\rceil}c_{1}(k,n)2^{\lfloor\frac{2n-k-3}{2}\rfloor}t^{n},$
(4.8) $U^{(1)}(y_{0}t^{k})=p_{1}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{0}(k,n)2^{\lfloor\frac{2n-k}{2}\rfloor}t^{n}+\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{1}(k,n)2^{\lfloor\frac{2n-k-4}{2}\rfloor}t^{n}.$
Proof.
The Appendix lists eight fundamental relations. The two fundamental relations of Group I fit the pattern of the relation (4.5) for two values of $k$. The same observations applies to the Groups II, III and IV with regard to the relations (4.6), (4.8), and (4.7), respectively. Applying Lemmas 4.1 and 4.2 immediately proves the statement for all $k\in\mathbb{N^{*}}$. ∎
We verify that
(4.9) $L_{1}=U^{(0)}(L_{0})=-4p_{0}t+y_{0}t,$
by Garvan’s MAPLE package ETA (3.1). Then we note that $p$ and $p_{1}$ have expansions in powers of $q$ with coefficients in $\mathbb{Z}$. Since $(1-q^{k})^{2}\equiv 1-q^{2k}\pmod{2}$, we have $\eta_{k}^{2}\equiv\eta_{2k}\pmod{2}$. Then
$2p=1+7\frac{\eta_{14}^{7}\eta_{4}^{5}\eta_{1}^{6}}{\eta_{28}^{3}\eta_{7}^{2}\eta_{2}^{13}}\equiv 1+7\frac{\eta_{14}^{7}\eta_{4}^{5}\eta_{2}^{3}}{\eta_{28}^{3}\eta_{14}\eta_{2}^{13}}\equiv 1+7\frac{\eta_{14}^{6}\eta_{4}^{5}}{\eta_{28}^{3}\eta_{2}^{10}}\equiv 8\equiv 0\pmod{2},$
which implies that function $p$ has expansions in powers of $q$ with coefficients in $\mathbb{Z}$. We now have that the coefficients of $4p_{1}=p+2pt+t+4t^{2}$ have integral coefficients. Hence, by the relations of $U^{(1)}(p_{0}t^{2})$ given in Appendix,
$U^{(1)}(p_{0}t^{2})=p_{1}t+4p_{1}f(t)+g(t),$
where $f(t)$ and $g(t)$ are polynomial of $t$ with integral coefficients, we see that $p_{1}t$ has expansions in powers of $q$ with coefficients in $\mathbb{Z}$ as well $p_{1}$. Furthermore, using (4.9) and combining the relations of $U^{(1)}(p_{0}t)$, $U^{(1)}(y_{0}t)$, $U^{(0)}(p_{1}t)$ given in Appendix and $U^{(0)}(t^{4})$ given in (3.3), we now state the following representation.
Lemma 4.5.
Given $L_{\alpha}$ in Definition 2.3, we have
(4.10) $L_{2}=$ $3p_{1}t-2t^{4}+4f(p_{1},t),$
$L_{3}=$ $y_{0}t+2y_{0}t^{2}+4g(p_{0},y_{0},t).$
for some polynomial $f$ and $g$ with integral coefficients.
Lemma 4.6.
For $\alpha\geq 1$ there exist integers $d_{n,i}^{(\alpha)}$, $i=0,1,2,3$ such that
(4.11) $L_{2\alpha-1}=p_{0}\sum_{k=1}^{\infty}d_{k,0}^{(2\alpha-1)}2^{k-1}t^{k}+y_{0}\sum_{k=1}^{\infty}d_{k,2}^{(2\alpha-1)}2^{k-1}t^{k},$
and
(4.12) $L_{2\alpha}=p_{1}\sum_{k=1}^{\infty}d_{k,1}^{(2\alpha)}2^{k-1}t^{k}+\sum_{k=3}^{\infty}d_{k,3}^{(2\alpha)}2^{k-3}t^{k}.$
Proof.
By (4.9) we see that $L_{1}=-4p_{0}t+y_{0}t$ has form given in (4.11). Assume that $L_{2\alpha-1}$ has the form given in (4.11) for a fixed $\alpha$, then
(4.13) $L_{2\alpha}=U^{(1)}(L_{2\alpha-1})=\sum_{k=1}^{\infty}d_{k,0}^{(2\alpha-1)}2^{k-1}U^{(1)}(p_{0}t^{k})+\sum_{k=1}^{\infty}d_{k,2}^{(2\alpha-1)}2^{k-1}U^{(1)}(y_{0}t^{k}).$
We aim to show that the form of each sum on the right-hand side of (4.13) satisfies the form in (4.12). From (4.7) and (4.8),
$L_{2\alpha}=$ $p_{1}\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{k,0}^{(2\alpha-1)}c_{0}(k,n)2^{k-1+\lfloor\frac{2n-k+1}{2}\rfloor}t^{n}+\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{k,0}^{(2\alpha-1)}c_{1}(k,n)2^{k-1+\lfloor\frac{2n-k-3}{2}\rfloor}t^{n}$
$+p_{1}\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{k,2}^{(2\alpha-1)}d_{0}(k,n)2^{k-1+\lfloor\frac{2n-k}{2}\rfloor}t^{n}+\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{k,2}^{(2\alpha-1)}d_{1}(k,n)2^{k-1+\lfloor\frac{2n-k-4}{2}\rfloor}t^{n}.$
For $k\geq 1$ and $n\geq\lceil\frac{k}{2}\rceil$ we have
$k-1+\lfloor\frac{2n-k+1}{2}\rfloor=\lfloor\frac{2n-k+1+2k-2}{2}\rfloor=n+\lfloor\frac{k-1}{2}\rfloor\geq n,$
$k-1+\lfloor\frac{2n-k}{2}\rfloor=\lfloor\frac{2n-k+2k-2}{2}\rfloor=n+\lfloor\frac{k-2}{2}\rfloor\geq n-1.$
For $k\geq 1$ and $n\geq\lceil\frac{k}{2}\rceil+2$ we have
$k-1+\lfloor\frac{2n-k-3}{2}\rfloor=\lfloor\frac{2n-k-3+2k-2}{2}\rfloor=n+\lfloor\frac{k-5}{2}\rfloor\geq n-2,$
$k-1+\lfloor\frac{2n-k-4}{2}\rfloor=\lfloor\frac{2n-k-4+2k-2}{2}\rfloor=n+\lfloor\frac{k-6}{2}\rfloor\geq n-3.$
So that the sums in the right-hand side of (4.13) has the form of (4.12). Hence $L_{2\alpha}$ has the desired form. The proof that the correct form of $L_{2\alpha}$ implies the correct form of $L_{2\alpha+1}$ is analogous. The general result follows by induction. ∎
By Lemma 4.2 and Lemma 4.6 we let
$L_{2\alpha-1}=p_{0}\sum_{n=1}^{\infty}l_{n,0}^{(2\alpha-1)}t^{n}+y_{0}\sum_{n=1}^{\infty}l_{n,2}^{(2\alpha-1)}t^{n},$
$L_{2\alpha}=p_{1}\sum_{n=1}^{\infty}l_{n,1}^{(2\alpha)}t^{n}+\sum_{n=3}^{\infty}l_{n,3}^{(2\alpha)}t^{n}.$
Next we define
$D^{(\alpha)}(l_{m,i},l_{n,j}):=l_{m,i}^{(\alpha)}l_{n,j}^{(\alpha+2)}-l_{m,i}^{(\alpha+2)}l_{n,j}^{(\alpha)},$
and denote $\pi(n)$ be the 2-adic order of n (i.e. the highest power of 2 that divides $n$). By Lemma 4.6, we should notice that for $m,n\in\mathbb{N}$ and $\alpha\geq 1$, $l_{n,1}$ and $l_{m,3}$ only appear in $L_{2\alpha}$, while $l_{n,0}$ and $l_{m,2}$ only appear in $L_{2\alpha-1}$.
Lemma 4.7.
For $\alpha\geq 1$ and $i,j\in\{0,1,2,3\}$, the following inequalities hold:
  • (1)
    for the odd cases we have
    (4.14) $\pi(D^{(2\alpha-1)}(l_{m,i},l_{n,j}))\geq 2\alpha-2+m+n+\lambda_{i,j},$
    (4.15) $\pi(D^{(2\alpha-1)}(l_{1,2},l_{2,0}))\geq 2\alpha,$
    (4.16) $\pi(D^{(2\alpha-1)}(l_{1,2},2l_{2,2}+l_{3,2}))\geq 2\alpha;$
  • (2)
    for the even cases we have
    (4.17) $\pi(D^{(2\alpha)}(l_{m,i},l_{n,j}))\geq 2\alpha-1+m+n+\lambda_{i,j}-\delta(m,i,n,j),$
    (4.18) $\pi(D^{(2\alpha)}(l_{1,1},l_{3,3}))\geq 2\alpha,$
    (4.19) $\pi(D^{(2\alpha)}(l_{2,1},l_{3,3}))\geq 2\alpha+1;$
where
$\lambda_{i,j}=\left\{\begin{array}[]{ll}-2,&\textrm{if $(i,j)=(1,1)$, $(0,2)$, $(2,0)$, or $(0,0)$},\\ -3,&\textrm{if $(i,j)=(2,2)$},\\ -4,&\textrm{if $(i,j)=(1,3)$ or $(3,1)$},\\ -6,&\textrm{if $(i,j)=(3,3)$},\end{array}\right.$
$\delta(m,i,n,j)=\left\{\begin{array}[]{ll}1,&\textrm{if $(m,i,n,j)\in G$ or $(n,j,m,i)\in G$},\\ 0,&\textrm{else},\end{array}\right.$
and
$G:=\{(m,i,n,j):(m,i)\in\{(5,1),(6,3),(7,3)\}~{}and~{}(n,j)\in\{(4,3),(5,3)\}\}.$
Proof.
We prove this lemma by mathematical induction. Firstly, we consider the case of $\alpha=1$. Since $l_{n,j}^{(1)}=0$ except $l_{1,0}^{(1)}=-4$ and $l_{1,2}^{(1)}=1$, from equation (4.11) we have
$\pi(D^{(1)}(l_{m,i},l_{n,j}))=\infty$
when $m,n\textgreater 1$,
$\pi(D^{(1)}(l_{1,i},l_{n,j}))=\pi(-D^{(1)}(l_{n,j},l_{1,i}))=\pi(l_{1,i}^{(1)}l_{n,j}^{(3)})\geq n+1+\lambda_{i,j}$
when $n\textgreater 1$, and
$\pi(D^{(1)}(l_{1,0},l_{1,2}))\geq 0=2+\lambda_{0,2},$
which prove (4.14) for $\alpha=1$. Specially, we obtain that
$\pi(D^{(1)}(l_{1,2},l_{2,0}))=\pi(l_{2,0}^{(3)})\geq 2,$
$\pi(D^{(1)}(l_{1,2},2l_{2,2}+l_{3,2}))=\pi(2l_{2,2}^{(3)}+l_{3,2}^{(3)})\geq 2,$
by Lemma 4.5, which prove that (4.15) and (4.16) hold for $\alpha=1$.
Secondly, we assume that (4.14)–(4.16) hold for $\alpha=s$ and prove that (4.17)–(4.19) hold for $\alpha=s$. We now compare the coefficients on both sides of
$L_{2s}=p_{1}\sum_{n=1}^{\infty}l_{n,1}^{(2s)}t^{n}+\sum_{n=3}^{\infty}l_{n,3}^{(2s)}t^{n}=\sum_{k=1}^{\infty}l_{k,0}^{(2s-1)}U^{(1)}(p_{0}t^{k})+\sum_{k=1}^{\infty}l_{k,2}^{(2s-1)}U^{(1)}(y_{0}t^{k}).$
Denote $p_{2}=y_{0}$ and $p_{3}=1$. For $u=0$ or $2$, define $x_{k,u}(m,i):=[p_{i}t^{m}]U^{(1)}(p_{u}t^{k})$ be the coefficient of $p_{i}t^{m}$ in $U^{(1)}(p_{u}t^{k})$, namely
$U^{(1)}(p_{u}t^{k})=p_{1}\sum_{n}x_{k,u}(n,1)t^{n}+\sum_{n}x_{k,u}(n,3)t^{n}.$
Then we derive that $l_{n,i}^{(2s)}=\sum\limits_{k,u}x_{k,u}(n,i)l_{k,u}^{(2s-1)}$ and
(4.20) $D^{(2s)}(l_{m,i},l_{n,j})=\sum_{k,u,r,v}x_{k,u}(m,i)x_{r,v}(n,j)D^{(2s-1)}(l_{k,u},l_{r,v}).$
By (4.7) and (4.8), we get $\pi(x_{k,u}(m,i))\geq\lfloor\frac{2m-k+\gamma_{u,i}}{2}\rfloor$, where $\gamma_{0,1}=1$, $\gamma_{2,1}=0$, $\gamma_{0,3}=-3$, and $\gamma_{2,3}=-4$. Now we first consider $(k,u,r,v)\in\{(1,2,2,2),(1,2,3,2),(3,2,1,2),(2,2,1,2)\}$. After checking all coefficients of $U^{(1)}(y_{0}t)$ and $U^{(1)}(y_{0}t^{2})$ given in the Appendix and $U^{(1)}(y_{0}t^{3})$ in (3.8), we obtain that
(4.21) $\pi(\sum_{r=2,3}{(x_{1,2}(m,i)x_{r,2}(n,j)+x_{r,2}(m,i)x_{1,2}(n,j))D^{(2s-1)}(l_{1,2},l_{r,2})})$
$\geq$ $2s-1+m+n+\lambda_{i,j}-\delta(m,i,n,j),$
which are based on inequalities (4.14) and (4.16) when $\alpha=s$.
Apart from $(k,u,r,v)\in\{(1,2,2,2),(1,2,3,2),(3,2,1,2),(2,2,1,2)\}$, we have
(4.22) $\pi(\sum_{k,u,r,v}x_{k,u}(m,i)x_{r,v}(n,j)D^{(2s-1)}(l_{k,u},l_{r,v}))$
$\geq$ $\min_{(k,u)\neq(r,v)}(k+r+2s-2+\lambda_{u,v}+\lfloor\frac{2m-k+\gamma_{u,i}}{2}\rfloor+\lfloor\frac{2n-r+\gamma_{v,j}}{2}\rfloor)$
$=$ $\min_{(k,u)\neq(r,v)}(m+n+2s-2+\lambda_{u,v}+\lfloor\frac{k+\gamma_{u,i}}{2}\rfloor+\lfloor\frac{r+\gamma_{v,j}}{2}\rfloor)$
$\geq$ $2s-1+m+n+\lambda_{i,j}.$
By (4.20)–(4.22) we get that inequality (4.17) holds for $\alpha=s$. For $\pi(D^{(2s)}(l_{1,1},l_{3,3}))$ and $\pi(D^{(2s)}(l_{2,1},l_{3,3}))$, based on Lemma 3.3 we derive:
(4.23) $l_{1,1}^{(2s)}=4l_{1,0}^{(2s-1)}+l_{2,0}^{(2s-1)}+35l_{1,2}^{(2s-1)}+7l_{2,2}^{(2s-1)},$
(4.24) $l_{2,1}^{(2s)}=40l_{1,0}^{(2s-1)}+20l_{2,0}^{(2s-1)}+6l_{3,0}^{(2s-1)}+l_{4,0}^{(2s-1)}+420l_{1,2}^{(2s-1)}+182l_{2,2}^{(2s-1)}+49l_{3,2}^{(2s-1)}+7l_{4,2}^{(2s-1)},$
(4.25) $l_{3,3}^{(2s)}=-4l_{1,0}^{(2s-1)}-l_{2,0}^{(2s-1)}-36l_{1,2}^{(2s-1)}-8l_{2,2}^{(2s-1)}.$
Rewriting $\pi(D^{(2s)}(l_{1,1},l_{3,3}))$ into $\pi(D^{(2s-1)}(4l_{1,0}+l_{2,0}+35l_{1,2}+7l_{2,2},-4l_{1,0}-l_{2,0}-36l_{1,2}-8l_{2,2}))$, we get $\pi(D^{(2s)}(l_{1,1},l_{3,3}))\geq 2s$ by (4.14) and (4.15). Similarly, we obtain $\pi(D^{(2s)}(l_{2,1},l_{3,3}))\geq 2s+1$ by (4.14).
Finally, we assume that (4.14)–(4.19) hold for $\alpha=s$ and prove that (4.14)–(4.16) hold for $\alpha=s+1$. For $u=1$ or $3$, define $y_{k,u}(m,i):=[p_{i}t^{m}]U^{(0)}(p_{u}t^{k})$ be the coefficient of $p_{i}t^{m}$ in $U^{(0)}(p_{u}t^{k})$, namely
$U^{(0)}(p_{u}t^{k})=p_{0}\sum_{n}y_{k,u}(n,0)t^{n}+y_{0}\sum_{n}y_{k,u}(n,2)t^{n}.$
Then
(4.26) $D^{(2s+1)}(l_{m,i},l_{n,j})=\sum_{k,u,r,v}y_{k,u}(m,i)y_{r,v}(n,j)D^{(2s)}(l_{k,u},l_{r,v}).$
With the help of (4.5) and (4.6), we have $\pi(y_{k,u}(m,i))\geq\lfloor\frac{2m-k+\gamma_{u,i}}{2}\rfloor$, where $\gamma_{1,0}=1$, $\gamma_{1,2}=-1$, $\gamma_{3,0}=2$, and $\gamma_{3,2}=2$. Apart from $(k,u,r,v)\in\{(1,1,3,3),(2,1,3,3),(3,3,1,1),(3,3,2,1)\}$, we obtain that
(4.27) $\pi(\sum_{k,u,r,v}y_{k,u}(m,i)y_{r,v}(n,j)D^{(2s)}(l_{k,u},l_{r,v}))$
$\geq$ $\min_{(k,u)\neq(r,v)}(m+n+2s-1+\lambda_{u,v}-\delta(k,u,r,v)+\lfloor\frac{k+\gamma_{u,i}}{2}\rfloor+\lfloor\frac{r+\gamma_{v,j}}{2}\rfloor)$
$\geq$ $2s+m+n+\lambda_{i,j},$
where the last inequality is determined by (4.17) when $\alpha=s$. By (4.18) and (4.19) we have
(4.28) $\pi(y_{1,1}(m,i)y_{3,3}(n,j)D^{(2s)}(l_{1,1},l_{3,3}))\geq 2s+m+n-2,$
(4.29) $\pi(y_{2,1}(m,i)y_{3,3}(n,j)D^{(2s)}(l_{2,1},l_{3,3}))\geq 2s+m+n-2.$
Hence (4.14) holds for $\alpha=s+1$ by (4.26)–(4.29). Next, we consider (4.15) and (4.16). By Lemma 3.3, we get that
(4.30) $l_{1,2}^{(2s+1)}=-l_{3,3}^{(2s)}-l_{1,1}^{(2s)},$
(4.31) $l_{2,2}^{(2s+1)}=-8l_{3,3}^{(2s)}-4l_{4,3}^{(2s)}-l_{5,3}^{(2s)}-10l_{1,1}^{(2s)}-5l_{2,1}^{(2s)}-l_{3,1}^{(2s)},$
(4.32) $l_{3,2}^{(2s+1)}=-28l_{3,3}^{(2s)}-32l_{4,3}^{(2s)}-18l_{5,3}^{(2s)}-6l_{6,3}^{(2s)}-l_{7,3}^{(2s)}-32l_{1,1}^{(2s)}-38l_{2,1}^{(2s)}-22l_{3,1}^{(2s)}-7l_{4,1}^{(2s)}-l_{5,1}^{(2s)},$
(4.33) $l_{2,0}^{(2s+1)}=64l_{3,3}^{(2s)}+28l_{4,3}^{(2s)}+7l_{5,3}^{(2s)}+84l_{1,1}^{(2s)}+38l_{2,1}^{(2s)}+8l_{3,1}^{(2s)}.$
Rewriting $\pi(D^{(2s+1)}(l_{1,2},l_{2,0}))$ by (4.30)–(4.33). Extracting terms which are less than $22s+$ 2 by (4.17) and (4.18), we get
$D^{(2s+1)}(l_{1,2},l_{2,0})\equiv D^{(2s)}(l_{3,3}+l_{1,1},l_{5,3}+2l_{2,1})\pmod{2^{2s+2}}.$
From (4.23)–(4.25) and
$l_{5,3}^{(2s)}=$ $-208l_{1,0}^{(2s-1)}-180l_{2,0}^{(2s-1)}-96l_{3,0}^{(2s-1)}-34l_{4,0}^{(2s-1)}-8l_{5,0}^{(2s-1)}-l_{6,0}^{(2s-1)}-2748l_{1,2}^{(2s-1)}$
$-1944l_{2,2}^{(2s-1)}-926l_{3,2}^{(2s-1)}-306l_{4,2}^{(2s-1)}-68l_{5,2}^{(2s-1)}-8l_{6,2}^{(2s-1)},$
which again derived by Lemma 3.3, we obtain that $\pi(D^{(2s)}(l_{3,3}+l_{1,1},l_{5,3}+2l_{2,1}))\geq 2s+2$ by $\pi(D^{(2s-1)}(l_{1,2},2l_{2,2}+l_{3,2}))\geq 2s$ and $\pi(D^{(2s-1)}(l_{1,2},l_{2,0}))\geq 2s$. Thus, $\pi(D^{(2s+1)}(l_{1,2},l_{2,0}))\geq 2s+2$, which means that (4.15) holds for $\alpha=s+1$. Similarly, rewriting $\pi(D^{(2s+1)}(l_{1,2},2l_{2,2}+l_{3,2}))$ by (4.30)–(4.32), we attain that each term is not less than $22s+$ 2 by (4.17) so that $\pi(D^{(2s+1)}(l_{1,2},2l_{2,2}+l_{3,2}))\geq 2s+2$. Thus, we prove that (4.15) and (4.16) hold for $\alpha=s+1$. ∎
Theorem 4.8.
For each $\alpha\geq 1$ there exist a constant $x_{\alpha}$ such that
$L_{2\alpha+2}\equiv x_{2\alpha}L_{2\alpha}\pmod{2^{2\alpha}},$
$L_{2\alpha+1}\equiv x_{2\alpha-1}L_{2\alpha-1}\pmod{2^{2\alpha-2}}.$
Proof.
By Lemma 4.7, we have
$\pi(D^{(2\alpha-1)}(l_{m,i},l_{n,j}))\geq 2\alpha-2,$
$\pi(D^{(2\alpha)}(l_{m,i},l_{n,j}))\geq 2\alpha.$
According to $U^{(1)}(y_{0}t)\equiv p_{1}t\pmod{2}$ and $U^{(0)}(p_{1}t)\equiv y_{0}t\pmod{2}$, we have
(4.34) $U^{(1)}(U^{(0)}(p_{1}t))\equiv p_{1}t\pmod{2},~{}~{}~{}\quad~{}~{}~{}U^{(0)}(U^{(1)}(y_{0}t))\equiv y_{0}t\pmod{2}.$
Using (4.9) and (4.10), we have $L_{1}\equiv l_{1,2}^{(1)}y_{0}t\equiv y_{0}t\pmod{2}$ and $L_{2}\equiv l_{1,1}^{(2)}p_{1}t\equiv p_{1}t\pmod{2}$, where $l_{1,2}^{(1)}=1$ and $2\nmid l_{1,1}^{(2)}$. Then, by (4.34), we have for $\alpha\geq 1$,
$L_{2\alpha-1}\equiv l_{1,2}^{(2\alpha-1)}y_{0}t\pmod{2},~{}~{}~{}\quad~{}~{}~{}L_{2\alpha}\equiv l_{1,1}^{(2\alpha)}p_{1}t\pmod{2}.$
Note that $2\nmid l_{1,1}^{(2\alpha)}$ and $2\nmid l_{1,2}^{(2\alpha-1)}$. Let $x_{2\alpha-1}$ be a solution of $l_{1,2}^{(2\alpha-1)}\equiv x_{2\alpha-1}l_{1,2}^{(2\alpha+1)}\pmod{2^{2\alpha-2}}$, then for $(n,i)\neq(1,2)$,
$l_{n,i}^{(2\alpha-1)}l_{1,2}^{(2\alpha-1)}\equiv l_{n,i}^{(2\alpha-1)}x_{2\alpha-1}l_{1,2}^{(2\alpha+1)}\equiv x_{2\alpha-1}l_{1,2}^{(2\alpha-1)}l_{n,i}^{(2\alpha+1)}\pmod{2^{2\alpha-2}}.$
Cancelling $l_{1,2}^{(2\alpha-1)}$, we obtain $l_{n,i}^{(2\alpha-1)}\equiv x_{2\alpha-1}l_{n,i}^{(2\alpha+1)}\pmod{2^{2\alpha-2}}$. Similarly, let $x_{2\alpha}$ be a solution of $l_{1,1}^{(2\alpha)}\equiv x_{2\alpha}l_{1,1}^{(2\alpha+2)}\pmod{2^{2\alpha}}$, then for $(n,i)\neq(1,1)$,
$l_{n,i}^{(2\alpha)}l_{1,1}^{(2\alpha)}\equiv l_{n,i}^{(2\alpha)}x_{2\alpha}l_{1,1}^{(2\alpha+2)}\equiv x_{2\alpha}l_{1,1}^{(2\alpha)}l_{n,i}^{(2\alpha+2)}\pmod{2^{2\alpha}}.$
Cancelling $l_{1,1}^{(2\alpha)}$, we obtain $l_{n,i}^{(2\alpha)}\equiv x_{2\alpha}l_{n,i}^{(2\alpha+2)}\pmod{2^{2\alpha}}$. Thus we complete the proof of Theorem 4.8. ∎
Based on Lemma 2.4 and Theorem 4.8, we can directly obtain the theorem below.
Theorem 4.9.
For $\alpha\geq 1$ and $n\geq 0$, we have
$\Delta_{3}(\lambda_{\alpha})\Delta_{3}(2^{\alpha+2}n+\lambda_{\alpha+2})\equiv\Delta_{3}(\lambda_{\alpha+2})\Delta_{3}(2^{\alpha}n+\lambda_{\alpha})\pmod{2^{2\lfloor\frac{\alpha}{2}\rfloor}},$
where
$\lambda_{\alpha}=\left\{\begin{array}[]{ll}\frac{2^{\alpha+1}+1}{3},{}&\text{if}~{}\alpha~{}\text{is even,}\\ \frac{2^{\alpha}+1}{3},{}&\text{if}~{}\alpha~{}\text{is odd.}\end{array}\right.$
Appendix. The fundamental relations
Group I:
$U^{(0)}(t)=$ $p_{0}(2^{4}t^{3}+3\cdot 2^{3}t^{2}+11\cdot 2t+7)+y_{0}(-2t-1);$
$U^{(0)}(t^{2})=$ $p_{0}(2^{6}t^{5}+2^{7}t^{4}+9\cdot 2^{4}t^{3}+19\cdot 2^{2}t^{2}+7\cdot 2t)+y_{0}(-2^{3}t^{3}-2^{3}t^{2}-2t).$
Group II:
$U^{(0)}(p_{1}t)=$ $p_{0}(2^{8}t^{7}+5\cdot 2^{7}t^{6}+7\cdot 2^{7}t^{5}+45\cdot 2^{4}t^{4}+21\cdot 2^{4}t^{3}+21\cdot 2^{2}t^{2}+2^{3}t)$
$+y_{0}(-2^{5}t^{5}-3\cdot 2^{4}t^{4}-2^{5}t^{3}-5\cdot 2t^{2}-t);$
$U^{(0)}(p_{1}t^{2})=$ $p_{0}(2^{10}t^{9}+3\cdot 2^{10}t^{8}+39\cdot 2^{7}t^{7}+39\cdot 2^{7}t^{6}+201\cdot 2^{4}t^{5}+167\cdot 2^{3}t^{4}+83\cdot 2^{2}t^{3}$
$+19\cdot 2t^{2})+y_{0}(-2^{7}t^{7}-2^{8}t^{6}-15\cdot 2^{4}t^{5}-2^{7}t^{4}-19\cdot 2t^{3}-5t^{2}).$
Group III:
$U^{(1)}(y_{0}t)=$ $p_{1}(2^{13}t^{9}+3\cdot 2^{13}t^{8}+69\cdot 2^{9}t^{7}+125\cdot 2^{8}t^{6}+617\cdot 2^{5}t^{5}+131\cdot 2^{6}t^{4}+597\cdot 2^{2}t^{3}$
$+105\cdot 2^{2}t^{2}+35t)-2^{13}t^{11}-7\cdot 2^{12}t^{10}-81\cdot 2^{9}t^{9}-75\cdot 2^{9}t^{8}-737\cdot 2^{5}t^{7}$
$-621\cdot 2^{4}t^{6}-687\cdot 2^{2}t^{5}-225\cdot 2t^{4}-9\cdot 2^{2}t^{3};$
$U^{(1)}(y_{0}t^{2})=$ $p_{1}(2^{15}t^{11}+7\cdot 2^{14}t^{10}+95\cdot 2^{11}t^{9}+103\cdot 2^{11}t^{8}+1251\cdot 2^{7}t^{7}+1371\cdot 2^{6}t^{6}+543\cdot 2^{6}t^{5}$
$+1211\cdot 2^{3}t^{4}+443\cdot 2^{2}t^{3}+91\cdot 2t^{2}+7t)-2^{15}t^{13}-2^{17}t^{12}-111\cdot 2^{11}t^{11}$
$-245\cdot 2^{10}t^{10}-1495\cdot 2^{7}t^{9}-817\cdot 2^{7}t^{8}-5\cdot 2^{13}t^{7}-1387\cdot 2^{3}t^{6}-243\cdot 2^{3}t^{5}-93\cdot 2t^{4}$
$-2^{3}t^{3}.$
Group IV:
$U^{(1)}(p_{0}t)=$ $p_{1}(2^{8}t^{7}+5\cdot 2^{7}t^{6}+11\cdot 2^{6}t^{5}+7\cdot 2^{6}t^{4}+11\cdot 2^{4}t^{3}+5\cdot 2^{3}t^{2}+2^{2}t)$
$-2^{8}t^{9}-3\cdot 2^{8}t^{8}-13\cdot 2^{6}t^{7}-17\cdot 2^{5}t^{6}-13\cdot 2^{4}t^{5}-11\cdot 2^{2}t^{4}-2^{2}t^{3};$
$U^{(1)}(p_{0}t^{2})=$ $p_{1}(2^{10}t^{9}+3\cdot 2^{10}t^{8}+33\cdot 2^{7}t^{7}+55\cdot 2^{6}t^{6}+121\cdot 2^{4}t^{5}+11\cdot 2^{6}t^{4}+5\cdot 2^{5}t^{3}$
$+5\cdot 2^{2}t^{2}+t)-2^{10}t^{11}-7\cdot 2^{9}t^{10}-39\cdot 2^{7}t^{9}-33\cdot 2^{7}t^{8}-145\cdot 2^{4}t^{7}$
$-103\cdot 2^{3}t^{6}-45\cdot 2^{2}t^{5}-5\cdot 2^{2}t^{4}-t^{3}.$
Acknowledgements